A psychologist believes that it will take at least an hour for certain disturbed children to learn a task. A

random sample of 35 of these children results in a mean of 50 minutes to learn the task. Should the

psychologist modify her belief at 0.01 level if the population standard deviation can be assumed to be 15

minutes.


1
Expert's answer
2021-06-30T06:45:13-0400

Solution:


Hypothesized Population Mean "\mu=60\ min"

Standard Deviation "\sigma=15\ min"

Sample Size "n=30"

Sample Mean "\bar{x}=50\ min"

Significance Level "\alpha=0.01"


Null and Alternative Hypotheses

The following null and alternative hypotheses need to be tested:

"H_0: \mu\geq60"

"H_1: \mu<60"

This corresponds to left-tailed test, for which a z-test for one mean, with known population standard deviation will be used.

Based on the information provided, the significance level is "\alpha=0.01,"and the critical value for left-tailed test is "z_c=-2.3263."

The rejection region for this left-tailed test is "R=\{t:z<-2.3263\}."


The "z" - statistic is computed as follows:



"z=\dfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\dfrac{50-60}{15/\sqrt{30}}\approx-3.6515"

Since it is observed that "z=-3.6515<-2.3263=z_c," it is then concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than "60," at the "\alpha=0.01" significance level.


Using the P-value approach: The p-value for left-tailed, the significance level "\alpha=0.01, z=-3.6515," is "p=0.00013," and since "p=0.00013<0.01=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is less than "60," at the "\alpha=0.01" significance level.